OpenStudy (donblue22):

Pleae help me simplify this algebraic fraction (not showing up right equation in box below)

OpenStudy (donblue22):

\[\frac{b^2-3b-10}{ (b-2)^2 }*\frac{ b-2 }{ b-5}\] is the algebraic fraction

OpenStudy (shadowlegendx):

Do they tell you what b is or do we have to find it?

OpenStudy (donblue22):

I dont have to solve only simplify

OpenStudy (cwrw238):

first factor the trinomial b^2 - 3b - 10 can you do that?

OpenStudy (cwrw238):

you need to numbers whose product is -10 and whose sum = -3 and they will be of the form (x + a)(x - b)

OpenStudy (donblue22):

(b+2)(b-5)

OpenStudy (cwrw238):

right now you can simplify

OpenStudy (donblue22):

Thanks

OpenStudy (cwrw238):

not that (b-5) and (b-2) are common factors in top and bottom of the fractions

OpenStudy (donblue22):

Yep I canceled out and got \[\frac{ b+2 }{ b-2 }\]